# Image-to-Video Transfer Learning based on Image-Language Foundation Models: A Comprehensive Survey

### v1

@@ Abstract
Image–Language Foundation Models (ILFM) have demonstrated remarkable success in image–text understanding/generation tasks, providing transferable multimodal representations that generalize across diverse downstream image-based tasks. The advancement of video-text research has spurred growing interest in extending image-based models to the video domain. This paradigm, known as image-to-video transfer learning, succeeds in alleviating the substantial data and computational requirements associated with training video-language foundation models from scratch for video-text learning. This survey provides the first comprehensive review of this emerging field, which begins by summarizing the widely used ILFM and their capabilities. We then systematically classify existing image-to-video transfer learning strategies into two categories: frozen features and modified features, depending on whether the original representations from ILFM are preserved or undergo modifications. Building upon the task-specific nature of image-to-video transfer, this survey methodically elaborates these strategies and details their applications across a spectrum of video-text learning tasks, ranging from fine-grained (e.g., spatio-temporal video grounding) to coarse-grained (e.g., video question answering). We further present a detailed experimental analysis to investigate the efficacy of different image-to-video transfer learning paradigms on a range of downstream video understanding tasks. Finally, we identify prevailing challenges and highlight promising directions for future research. By offering a comprehensive and structured overview, this survey aims to establish a structured roadmap for advancing video-text learning based on existing ILFM, and to inspire future research directions in this rapidly evolving domain.
@@
Foundation Models, Video Understanding, Transfer Learning, Vision-Language, Image-to-Video Adaption.
@@ I Introduction
Multimodal learning, particularly vision-language (or vision-text) learning, has attracted significant attention in the community in recent years. With larger datasets and more powerful computing resources, researchers can train image-language foundation models that can adapt to many downstream tasks and modalities. Image-based multi-modality systems such as CLIP[1] and BLIP[2] showed that training models with pairs of images and texts at scale can yield widely transferable features. Newer models like MDETR[3] and GroundingDINO[4] go further by aligning free-form language with locally aligned image regions, closing the gap between textual description and image location. Building on these successes, researchers are actively exploring how to migrate these models, proven successful on static images, to the more complex video domain. The goal is to address the unique challenges posed by the additional temporal dimension in videos and to leverage the rich dynamic information.
@@
Figure 1: Comparison of two representative strategies for addressing video-text understanding problem with image-text and video-text foundation models. Pretraining video-based foundation models is much more challenging than pretraining image-based foundation models, as they often contain more parameters and require larger amounts of training data and computational resources.
@@
Over recent years, image-to-video transfer learning, which adapts knowledge from vision-language foundation models to the video domain, has emerged as a prominent paradigm for video-language understanding. Alternatively, another approach involves pre-training video-language foundation models directly from massive video-text data before transferring them to downstream tasks. As shown in Fig 1, compared to direct pre-training—which faces challenges such as high data collection costs and low computational efficiency—we find that initializing from large-scale image-text pre-trained models offers greater practicality and scalability. This is attributed to the greater abundance of image-text data and lower associated costs. For instance, UniFormerV1 [5] was pre-trained directly on the Kinetics-400 [6] dataset, whereas UniFormerV2 [7] underwent post-pre-training starting from an image-text model. Compared to V1, UniFormerV2 achieves superior performance with reduced training overhead, demonstrating that adapting image representations aligned with text to various video scenarios via task-specific transfer techniques is both more flexible and cost-effective.
@@
However, extending the notable success of image-language foundation models to the video domain presents significant challenges, primarily arising from the temporal dimension. In contrast to static images, videos comprise sequences of frames that encode complex spatiotemporal dynamics, including motion, event progression, and causal interactions. Consequently, a fundamental architectural paradigm shift is necessary, moving from processing static spatial information to effectively modeling these intricate spatial-temporal relationships. Early methods processed each frame with 2D image networks and then applied a simple temporal fusion layer, but this setup cannot capture fine motion cues. Recently, researchers are adapting these models to video domain by inserting lightweight temporal modules into a frozen backbone and applying parameter-efficient fine-tuning techniques like adapters, often employing these strategies in hybrid configurations. Furthermore, a growing number of diverse approaches are continuously being introduced and explored in this domain. Side-tuning approaches [8, 9], for example, is becoming increasingly popular as one of the most promising ways to accommodate the critical challenge of image-to-video transfer learning brought by the huge parameter and computation overheads. Other methods like distillation, also shows further potential to adapt the image-text aligned knowledge into the video domain especially under the era of Multimodal Large Language Models (MLLMs) in which we can capitalize on the exceptionally strong multimodal knowledge contained by these excessively large models.
@@
Despite the existence of numerous paradigms for image-to-video transfer learning, it remains underexplored which paradigm or combination thereof is most suitable for specific downstream tasks, as different tasks require models to possess distinct comprehension capabilities. For example, spatio-temporal video grounding (STVG) requires a model to enhance its capability for fine-grained temporal event localization while preserving its existing static spatial grounding performance. This objective is well-suited for the integration of supplementary video-level features. In contrast, Video Question Answering (Video QA) necessitates only a coarse-grained comprehension of video events but demands a more robust capacity for multi-modal knowledge understanding. The optimal transfer learning paradigm for this latter task remains an open question.
@@
In this survey, we provide a comprehensive overview of established strategies for migrating image foundation models to various video-centric tasks. Beyond summarizing these techniques, we offer a critical analysis of their suitability for different application scenarios, concluding with a discussion on the prevailing challenges and promising directions for future research. The rest of this survey is organized as follows. Section II introduce the commonly used vision-language foundation models developed in the community. Section III reviews and categorizes the techniques of leveraging image-level knowledge from Image-Language Foundation Models (ILFM) for video-language understanding. Section IV and V systemically review the image-to-video transfer learning approaches for addressing varied downstream video-text understanding tasks including fine-grained to coarse-grained problems, respectively. In Section VI, we provide the detailed experimental analysis. Finally, we provide several promising image-to-video transfer learning research directions in Section VII and conclude in Section VIII.
II Overview of Vision-Language Foundation Model
Pretrained ILFM that leverage both visual and textual information have revolutionized many visual perception tasks including object detection [10, 11, 12], image captioning [13, 14, 15], and visual question answering (VQA)[16, 17, 18], etc. Among the most influential models in this domain, several have introduced foundational innovations: CLIP[1] establishes image-text alignment through contrastive pretraining; BLIP[2] and BLIP-2[19] improve vision-language representation learning with bootstrapped captioning and parameter-efficient architecture; MDETR[3] and GroundingDINO[4] advance open-vocabulary detection and referring comprehension using language-aware visual encoding; and LLaVA[20] integrates visual encoders with large language models to support instruction-based multimodal dialogue. Collectively, these models bridge the gap between vision and language understanding and provide a diverse set of powerful and adaptable foundations for a wide spectrum of downstream vision-language applications.
@@ II Overview of Vision-Language Foundation Model
TABLE I: Classification and Analysis of Image-Language Foundational Models
@@
Foundation Model Core Architecture Pre-training Objective(s) Pre-training Data Primary Output Key Capability
@@
CLIP[1] Dual-Encoder (Image Encoder + Text Encoder) Image-Text Contrastive Learning 400M image-text pairs (Web-crawled) Aligned Image & Text Embeddings Zero-Shot Image Classification, Robust Feature Representation
BLIP & BLIP2[2, 19] Multimodal Mixture of Encoder-Decoder (ViT, Q-Former, Text Encoder/Decoder) Contrastive, Matching, Captioning Approx. 14M image-text pairs (Academic datasets + Web data) Embeddings, Matching Scores, Text Visual Question Answering (VQA), Image Captioning
@@
MDETR[3] Transformer Encoder-Decoder Phrase-to-Box Grounding 1.3M images with 11.5M phrase-box annotations (Flickr30k, GQA, etc.) Bounding Boxes, Class Labels Open-Vocabulary Object Detection & Grounding
GroundingDINO[4] Transformer-based Detector with Image & Text Backbones Open-Set Object Detection Combination of COCO, Objects365, GoldG, etc. (Detection & Grounding datasets) Bounding Boxes, Class Labels Open-Set Detection
@@
LLaVA[20] Vision-Language Model(Image Encoder + Projection + LLM) Instruction Tuning on Image-Text Data Approx. 158K language-image instruction-following samples (GPT-4 generated) Textual Response Conversational VQA, Visual Reasoning
MDETR111ghttps://github.com/ashkamath/mdetr [3] is a foundational vision-language model designed for grounded visual-language understanding, where it predicts object locations in conjunction with their corresponding textual descriptions. As illustrated in Figure 2(a), the model employs dual encoders: a convolutional visual encoder (e.g., ResNet-101[21] or EfficientNet[22]) and a transformer-based text encoder (e.g., RoBERTa[23]). These encoders project both images and text into a shared latent space, where semantic alignment is enforced by maximizing similarity for matched pairs and minimizing it for mismatched ones. This joint representation allows MDETR to reason over spatial and semantic relationships across modalities, making it particularly effective for tasks such as phrase grounding, referring expression segmentation, and visual question answering (VQA).
MDETR is pre-trained on a curated corpus of approximately 200k images encompassing around 1.3 million region–phrase pairs.These aligned pairs are sourced from diverse datasets including Flickr30k[24], MS-COCO[25], and Visual Genome[26], ensuring dense grounding of objects to textual descriptions. The model employs a graph-coloring technique to consolidate multiple phrases per image, ensuring non-overlapping and semantically coherent region assignments that facilitate more effective representation learning. Trained at this scale, MDETR demonstrates strong zero-shot generalization capabilities across a variety of downstream tasks, including open-domain visual retrieval and fine-grained visual recognition.
@@
GroundingDINO [^2] https://github.com/IDEA-Research/GroundingDINO[4] introduces a novel approach to open-set visual grounding, designed to localize and identify objects based on free-form textual descriptions. As illustrated in Figure 2(b), the model builds upon a transformer-based architecture that integrates visual and linguistic modalities through a dedicated fusion module, enabling fine-grained cross-modal alignment between image regions and text queries. This design allows the model to effectively ground natural language phrases in images, making it highly suitable for tasks such as referring expression comprehension and open-vocabulary detection.The model is trained on a large-scale multi-source dataset comprising image-text pairs with bounding box annotations. Key data sources include Flickr30K Entities [24], MS-COCO [25], Visual Genome [26], and Objects365 [27], which provide diverse and fine-grained grounding supervision across common and long-tail categories. This extensive pretraining enables robust representation learning for both visual and textual modalities.
A major innovation of GroundingDINO lies in its integration of the DINO (Distilled Knowledge) framework into visual grounding. Through cross-modal knowledge distillation, the model captures fine-grained semantic information from textual descriptions and transfers this understanding to visual contexts with high precision. This mechanism enables GroundingDINO to localize objects in unseen images using only text prompts—without any task-specific retraining—showcasing remarkable zero-shot capability. This strength makes it particularly suitable for applications in dynamic visual environments and open-ended image retrieval tasks.
Figure 2: Popular image-textual foundation models developed in the community, which are widely used in the image-to-video transfer learning for addressing downstream video-language understanding problems.
CLIP (Contrastive Language-Image Pre-training) [^3] https://github.com/openai/CLIP [1] is an image-language foundational model that learns joint representations of images and text through contrastive pre-training. As Figure 2(c) illustrate, its architecture consists of dual encoders: a vision encoder built on scalable neural networks such as vision transformers or convolutional architectures, and a text encoder based on transformer networks. These encoders are trained to map paired images and texts into a shared latent space, where semantic alignment is achieved by maximizing similarity scores for correct image-text pairs and minimizing them for incorrect ones. The model is pre-trained on a dataset of 400 million publicly available image-text pairs collected from the web, enabling broad semantic understanding. A key innovation of CLIP is its zero-shot generalization capability, enabling it to be transferrable to other perception tasks, such as image classification and cross-modal retrieval, without requiring task-specific fine-tuning. Empirical studies show that CLIP performs competitively with fully supervised models across 30 diverse vision benchmarks, particularly excelling in fine-grained recognition and open-domain retrieval.
@@
   Frozen Features (§III-A)    Knowledge Distillation       (§III-A1) Fine-grained: OVTrack [28],OVTrack+ [29],VOVTrack [30],SLAck [31],GLATrack [32],UniVTG [33]. Coarse-grained: VideoDistill [34], Clip2tv [35], Text4Vis [36], FitCLIP [37], CLIPPING [38]. TeachCLIP [39],    Post-Network Tuning       (§III-A2) Fine-grained: NetTrack [40], OVTR [41], LaMOTer [42], OV2Seg [43], UnLoc [33], CLIP-VIS [44], MomentDETR [45], BAM-DETR [46], UniVTG [33], UnLoc [47], TFVTG [48], CoSPaL [49], STPro [50], BriVIS [51], OVFormer [52], DVIS++ [53]. Coarse-grained: BIMBA [54], LinVT [55], CLIP4Caption [56], E2DVC [57], HierarQ [58], UniFormerV2 [7].       Side-Tuning       (§III-A3) Fine-grained: R2-tuning [59], VDI [60], PATF [61], Tune-A-Video [62]. Coarse-grained: CLIP2Video [63], STAN [64], EVL [65], DiST [66], MAMS [67]. Modified Features (§III-B)       Full Fine-Tuning       (§III-B1) Fine-grained: STVGBERT [68], SWINBERT [69], TubeDETR [70], STCAT [71], ReferDINO [72], GEN1 [73]. Coarse-grained: Stable Video Diffusion [74], Stable Diffusion [75], I2VGen-XL [76], CLIP4Clip [77], CLIP-Hitchhiker [78], ViFi-CLIP [79], TVR [80], CenterCLIP [81].       Partial Tuning       (§III-B2) Fine-grained: Tune-A-Video [62], FRESCO [82]. Coarse-grained: VideoCrafter1 [83], TC-LLaVa [84], ModelScope Text-to-Video [85], TEA [86], mPLUG-2 [87]. Fine-Tuning with Extra Models       (§III-B3) Fine-grained: iKUN [88], CoSD [89], CoSTA [90], CGSTVG [91], TASTVG [92], Vtimellm [93], Seq2Time [94], NumPro [95], Chrono [96], NetTrack [40], InstFormer [97], Grounded-SAM-2 [98], AL-Ref-SAM-2 [99]. Coarse-grained: VTG-LLM [100], X-pool [101], MoMa [102], Vid2Seq [103], CM2 [104], EvCap [105], BIKE [106].    Fine-Tuning with Adapter       (§III-B4) Fine-grained: MASA [107], LlaViLo [108], ReVisionLLM [109], iKUN [88]. Coarse-grained: Tem-Adapter [110], ST-Adapter [111], Adaptformer [112], AIM [113], DUALPATH [114], TS2-Net [115], X-CLIP [116], D2ST [117], TR-Adapter [118].     Fine-Tuning with LoRA       (§III-B5) Fine-grained: ReferDINO [72], VTG-LLM [100], VTimeLLM [93], ReVisionLLM [109], NumPro [95], InstFormer [97]. Coarse-grained: DiscoVLA [119], HierarQ [58].       Prompt Tuning       (§III-B6) Coarse-grained: ATP [120], ActionCLIP [121], X-CLIP [116], ViTiS [122], Q-ViD [123], X-CLIP [124]. VOP [125], Clip-vip [126].
@@
Figure 3: Taxonomy of Image-to-video Transfer Learning Methods. Existing methods are categorized according to the employed transfer learning mechanisms, forming two major groups: frozen features and modified features. The frozen-feature group includes approaches such as knowledge distillation, post-network tuning, and side-tuning, while the modified-feature group encompasses various fine-tuning strategies, including full fine-tuning, adapter-based tuning, LoRA, and prompt tuning. Each category is further divided into fine-grained and coarse-grained paradigms, reflecting different levels of temporal and semantic alignment between visual and textual modalities.
BLIP [^4] https://github.com/salesforce/BLIP [2] (Bootstrapping Language-Image Pre-training) is a unified vision-language framework designed to jointly optimize understanding and generation tasks through multimodal co-training. As Figure 2(d) shows, its architecture integrates a vision encoder for image feature extraction, a text encoder-decoder for language processing, and a cross-modal transformer that dynamically fuses visual and textual representations. This hybrid design enables versatile applications, including visual question answering, image captioning, and text-guided image retrieval.
BLIP is effectively trained with a two-stage pre-training strategy. In the first stage, it employees a data bootstrapping mechanism to filter out noisy web-collected image-text pairs and generate high-quality synthetic captions via knowledge distillation. In the second stage, it optimizes the parameters by considering image-text contrastive learning to align modalities, image-text matching to verify cross-modal consistency, and generative tasks to produce contextually relevant textual outputs. Pre-trained on a combination of 129 million curated and 140 million web-sourced image-text pairs, BLIP achieves state-of-the-art results across 14 benchmarks, demonstrating strong robustness in low-data regimes.
BLIP-2555https://github.com/salesforce/LAVIS/tree/main/projects/blip2 [19] further improves BLIP by additionally introducing a lightweight querying module to bridge pre-trained vision encoders and large language models, in order to align visual features with textual embeddings. This delivers significant accuracy gains on some downstream visual-textual understanding tasks like visual question answering, with introducing considerable computational overhead. BLIP and its extensions BLIP-2 have been widely used in the community for addressing multimodal reasoning tasks including healthcare diagnostics and interactive recommendation systems.
LLaVA [^6] https://github.com/haotian-liu/LLaVA (Large Language and Vision Assistant) [20] is a vision-language model that integrates visual understanding with natural language processing through an end-to-end trainable architecture. As Figure 2(e) shows, The model consists of a pre-trained vision encoder CLIP-ViT-L/14 and a large language model (LLM) to process both textual and visual features. LLaVA employs a simple projection layer to align visual representations with the language model’s embedding space, enabling cross-modality fusion.
The training of LLaVA is split into two stages to ensure better cross-modality alignment and task-specific adaptability. In the first stage (i.e., feature alignment pretraining), the projection layers in LLaVA are learned with freezing vision encoder, which aims to learn a good semantic alignment between visual and textual representations. The pretraining is conducted on a subset of CC3M (Conceptual Captions 3M)[127], a dataset comprising 3 million image-text pairs. In the second stage (instruction fine-tuning), a large scale of instruction-following data generated by GPT-4, which includes approximately 158K multi-modal instruction-response pairs covering diverse tasks such as visual dialogue, detailed descriptions, and complex question answering are used to enhance the model’s reasoning and conversational capabilities. In this stage, both the projection layer and the LLM are updated and the vision encoder is fixed.

### v2

@@ Abstract
Image–Language Foundation Models (ILFMs) have demonstrated remarkable success in vision-language understanding, providing transferable multimodal representations that generalize across diverse downstream image-based tasks. The advancement of video-text research has spurred growing interest in extending image-based models to the video domain. This paradigm, termed as image-to-video transfer learning, effectively mitigates the substantial data and computational demands compared to training video-language models from scratch while achieves comparable or even stronger model performance. This survey provides the first comprehensive review of this emerging field, which begins by summarizing the widely used ILFMs and their capabilities. We then systematically classify existing image-to-video transfer learning techniques into two broad root categories (frozen features and adapted features), along with numerous fine-grained subcategories, based on the paradigm for transferring image understanding capability to video tasks. Building upon the task-specific nature of image-to-video transfer, this survey methodically elaborates these strategies and details their applications across a spectrum of video-text learning tasks, ranging from fine-grained settings (e.g., spatio-temporal video grounding) to coarse-grained ones (e.g., video question answering). We further present a detailed experimental analysis to investigate the efficacy of different image-to-video transfer learning paradigms on a range of downstream video understanding tasks. Finally, we identify prevailing challenges and highlight promising directions for future research. By offering a comprehensive and structured overview, this survey aims to establish a structured roadmap for advancing video-text learning based on existing ILFM, and to inspire future research directions in this rapidly evolving domain. Additional resources and updates are maintained at our real-time github repository: https://github.com/YuriPreisdent/awesome-image-to-video-transfer.
@@
Foundation Models, Video Understanding, Transfer Learning, Vision-Language, Image-to-Video Transfer.
@@ I Introduction
Multimodal learning, particularly vision-language (or image-text) learning, has attracted significant attention in the community in recent years. With larger datasets and more powerful computing resources, researchers can train ILFMs to adapt to various downstream tasks and modalities. Vision-language models such as CLIP [146] and BLIP [90] demonstrate that large-scale pre-training on image-text pairs yields highly transferable visual representations. Subsequent approaches including MDETR [73] and Grounding-DINO (G-DINO) [117] advance this paradigm by explicitly aligning free-form linguistic expressions with localized image regions, thereby bridging the gap between open-vocabulary language and visual grounding. More recently, LLM (Large Language Model)-based ILFMs[114, 3, 25] have emerged, which integrate diverse image-text understanding capabilities and can generate contextually appropriate responses in accordance with human instructions. Given that a video is fundamentally a sequence of images, a key research direction involves transferring the capabilities of ILFMs, which have demonstrated effectiveness on static image understanding, to the more complex video domain for solving challenging video understanding problems.
@@
Figure 1: Comparison of two representative strategies for addressing video-text understanding problem with image-text and video-text foundation models. Pretraining video-based foundation models is much more challenging or even impractical, as they often contain more parameters and require larger amounts of training data and computational resources.
Figure 2: A comprehensive timeline depicting the development of image-language foundation models over recent years.
@@
Compared to directly transferring models pre-trained solely on large-scale video-text data, adapting ILFMs to downstream video tasks demonstrates greater advantages. As illustrated in Fig. 1, training video-language foundation models from scratch encounters key challenges such as low computational efficiency and high data collection costs, since acquiring high-quality video-text data is inherently extremely difficult. In contrast, initializing from large-scale image-text pre-trained models presents higher practicality and scalability, owing to the richer availability and lower cost of image-text data. For instance, UniFormerV1 [93] was pre-trained directly on the Kinetics-400 [76] dataset, whereas UniFormerV2 [92] underwent post-pre-training starting from an image-text model. Compared to V1, UniFormerV2 achieves superior performance across diverse video understanding tasks with reduced training overhead, demonstrating that adapting image representations aligned with text to various video scenarios via task-specific transfer techniques is both more flexible and cost-effective.
@@
The primary challenge in extending ILFMs to the video domain stems from the effective modeling of temporal information. In contrast to static images, videos comprise sequences of frames that encode complex spatiotemporal dynamics, including motion, event progression, and causal interactions. Consequently, an architectural transferring paradigm is necessary, moving from processing static spatial information to effectively modeling these intricate spatial-temporal relationships. Early methods processed each frame with 2D image networks and then applied a simple temporal fusion layer, which cannot effectively capture fine motion cues. Recently, researchers have been adapting these models to the video domain by inserting lightweight temporal modules into a frozen backbone and applying parameter-efficient fine-tuning techniques like adapters, often employing these strategies in hybrid configurations. Furthermore, a growing number of diverse approaches are continuously being introduced and explored in this domain. Side-tuning approaches [204, 157], for example, is becoming increasingly popular as one of the most promising ways to accommodate the critical challenge of image-to-video transfer learning brought by the huge parameter and computation overheads. Other methods like distillation, also shows further potential to adapt the image-text aligned knowledge into the video domain especially under the era of Multimodal Large Language Models (MLLMs) in which we can capitalize on the exceptionally strong multimodal knowledge contained by these excessively large image-based MLLMs.
@@
Despite the existence of numerous paradigms for image-to-video transfer learning, it remains underexplored which paradigm or combination thereof is most suitable for specific downstream tasks, as different tasks require models to possess distinct comprehension capabilities. For example, Spatio-Temporal Video Grounding (STVG) requires a model to enhance its capability for fine-grained temporal event localization while preserving its existing static spatial grounding performance. This objective is well-suited for the paradigm of fine-tuning with extra video models. In contrast, Video Question Answering (VideoQA) necessitates only a coarse-grained comprehension of video events but demands a more robust capacity for multi-modal knowledge understanding. Up to now, the optimal transfer learning paradigm for this latter task still remains an open question in the community.
@@
In this survey, we provide a comprehensive overview of established strategies for migrating ILFMs to video-centric tasks. Beyond summarizing these techniques, we offer a critical analysis of their suitability for different application scenarios, concluding with a discussion on the prevailing challenges and promising directions for future research. The rest of this survey is organized as follows. Section II introduces the commonly used ILFMs developed in the community. Section III reviews and categorizes the techniques of leveraging image-level knowledge from ILFMs for video-language understanding. Sections IV and V systemically review the image-to-video transfer learning approaches for addressing varied downstream video-text understanding tasks, including fine-grained and coarse-grained problems, respectively. In Section VI, we provide the detailed experimental analysis. Finally, we provide promising image-to-video transfer learning research directions in Section VII and conclude in Section VIII.
II Overview of Image-Language
Foundation Models
@@ II Overview of Image-Language
TABLE I: Categorization and Analysis of popular Image-Language Foundational Models (ILFMs)
@@
Foundation Model Pre-training Objective(s) Pre-training Data Primary Output Key Capability
@@
CLIP [146] Image-Text Contrastive Learning 400M image-text pairs Aligned Image & Text Embeddings Zero-Shot Image Classification, Robust Image-text Representation
@@
MDETR [73] Phrase-to-Box Grounding 1.3M images with 11.5M phrase-box annotations Bounding Boxes, Class Labels Open-Vocabulary Object Detection & Grounding
G-DINO [117] Open-Set Object Detection Over 3M images with 30M+ box annotations and 5M+ phrase relations Bounding Boxes, Class Labels Open-Vocabulary Object Detection & Grounding
@@
LLaVA [114] Instruction Tuning on Image-Text Data 158K image-text instruction samples Textual Response Conversational VQA, Visual Reasoning
InternVL [25] Image–Text Pre-training, Generative Training, Instruction Tuning 4B+ image–text pairs & 100M+ curation data & 1.5M+ instruction data Aligned Vision–Language Embeddings, Textual Responses Generic Visual Recognition, Image Retrieval, Captioning, VQA, Multi-Modal Dialogue
QwenVL [3] Image–Text Pretraining, Multi-Task Learning, Instruction Tuning 1.4B+ image-text pairs & Curated Billion-scale Bilingual Corpus Textual Responses, Localized Bounding Boxes Image Captioning, VQA, OCR-based Text Reading, Visual Grounding, Multilingual Dialogue
Pretrained ILFMs have revolutionized a broad spectrum of visual understanding tasks such as object detection [51, 52, 149], image captioning [192, 199, 72], visual question answering [41, 42, 12], etc. In this section, our goal is to introduce several representative ILFMs: CLIP [146] establishes image-text alignment through contrastive pretraining; MDETR [73] and G-DINO [117] advance open-vocabulary detection and referring comprehension using language-aware visual encoding; lately LLaVA [114], InternVL [25] and Qwen-VL [3] integrate visual encoders with capable LLMs for instruction-based multimodal understanding. Collectively, these models bridge the gap between vision and language, providing powerful and adaptable foundations for various downstream vision-language applications.
MDETR [73] is a representative foundation model for end-to-end grounded multimodal understanding, which can detect object locations in accordance with their textual descriptions in the given language query. As illustrated in Figure 3(a), it first employs a convolutional visual encoder (e.g., ResNet-101 [60]) and a transformer textual encoder [119] to project both images and texts into a joint embedding space. Then a cross-encoder and a transformer decoder are utilized in cascade to obtain a grounded multimodal representation, on top of which a fine-grained contrastive loss aligns words or phrases in the text with their matching objects in the image. Learning such a grounded representation enables MDETR to reason over spatial relations and semantic bindings across modalities. Furthermore, MDETR is pre-trained on a massive data scale of $\mathord{\sim}$1.3M region-phrase pairs, which ensures robust learning of dense grounding between image objects and text tokens. Built upon the large-scale pretraining, MDETR exhibits strong generalization performance across various fine-grained tasks such as phrase grounding and referring expression comprehension.
@@
G-DINO [117] aims to localize and identify objects based on free-form textual descriptions. As illustrated in Figure 3(b), the model builds upon a transformer-based architecture that integrates visual and linguistic modalities through a dedicated fusion module, enabling fine-grained cross-modal alignment between image regions and text queries. This design allows the model to effectively ground natural language phrases in images. The model is trained on a large-scale multi-source dataset comprising image-text pairs with bounding box annotations, which enables G-DINO to localize objects in unseen images using only text prompts—without any task-specific retraining, showcasing remarkable zero-shot capability. This strength makes it particularly suitable for applications in dynamic visual environments and open-ended retrieval tasks.
CLIP [146] is one of the earliest and most popular ILFMs that learns aligned representations of images and texts via large-scale contrastive pretraining. As illustrated in Figure 3(c), CLIP employs a dual encoder framework: a vision encoder built on scalable architectures such as vision transformers or convolutional networks, and a text encoder based on transformers. These encoders are trained to map paired images and texts into a shared latent space, where semantic alignment is achieved by maximizing similarity scores for relevant image-text pairs and minimizing them for irrelevant ones. A large-scale dataset consisting of $\mathord{\sim}$400M image-text pairs is collected from the Internet to train CLIP, enabling general vision-language alignment. Building upon this, subsequent research has focused on developing improved training recipes for CLIP-like pretraining for better scalability and training stability. Notably, OpenCLIP [67] introduces an open-source framework to reproduce and extend the training of CLIP model. SigLIP [202] is proposed to replace the standard InfoNCE objective with a simpler pairwise sigmoid loss. More recently, SigLIP2 [165] makes further extensions by additionally including captioning-based objectives and self-supervised losses.
Figure 3: Popular ILFMs for image-to-video transfer learning in video-language understanding tasks.
@@
   Frozen Features (§III-A)      Knowledge Distillation        (§III-A1) Fine-grained: OVTrack [97], OVTrack+ [103], VOVTrack [144], OVSORT [95], SLAck [99], GLATrack [86], UniVTG [109], Coarse-grained: VideoDistill [219], CLIP2TV [47], Text4Vis [183], MS-SL [32], FitCLIP [13], CLIPPING [141], TeachCLIP [163], MS-SL++ [23].       Post-Network Tuning        (§III-A2) Fine-grained: NetTrack [212], OVTR [88], LaMOTer [101], OV2Seg [169], UnLoc [109], CLIP-VIS [217], MomentDETR [85], BAM-DETR [83], UniVTG [109], UnLoc [194], TFVTG [213], CoSPaL [81], STPro [48], BriVIS [26], OVFormer [39], SpaceVLLM [170], VideoTG-R1 [33], DVIS++ [206]. Coarse-grained: BIMBA [68], LinVT [45], CLIP4Caption [161], E2DVC [182], HierarQ [2], LeAdQA [34], UniFormerV2 [92], SG-VLM [127], OHRC [176], TubeRMC [87].         Side-Tuning        (§III-A3) Fine-grained: R2-tuning [118], VDI [125], PATF [134]. Coarse-grained: CLIP2Video [38], STAN [115], EVL [111], DiST [145], MoTED [207], MAMS [84]. Adapted Features (§III-B)        Full Fine-Tuning        (§III-B1) Fine-grained: STVGBERT [155], SWINBERT [108], TubeDETR [195], STCAT [71], ReferDINO [104], Coarse-grained: CLIP4Clip [126], CLIP-Hitchhiker [5], ViFi-CLIP [147], TVR [173], CenterCLIP [211], X-CLIP [129].         Partial Tuning        (§III-B2) Coarse-grained: Token-Mixing [121], TC-LLaVa [46], TEA [208], mPLUG-2 [189]. Fine-Tuning with Extra Models        (§III-B3) Fine-grained: iKUN [36], CoSD [110], CoSTA [105], CGSTVG [55], TASTVG [56], Vtimellm [63], Seq2Time [30], NumPro [185], Chrono [131], NetTrack [212], InstFormer [57], Grounded-SAM-2 [150], UniTime [102], AL-Ref-SAM-2 [64]. Coarse-grained: VTG-LLM [58], X-Pool [53], MoMa [198], Vid2Seq [196], CM2 [77], EvCap [116], BIKE [184].     Fine-Tuning with Adapter        (§III-B4) Fine-grained: MASA [98], LlaViLo [128], ReVisionLLM [59], iKUN [36]. Coarse-grained: Tem-Adapter [16], ST-Adapter [137], Adaptformer [22], AIM [197], DUALPATH [140], TS2-Net [120], D2ST [142], MV-Adapter [70], Q-adapter [20], ZeroI2V [100], TR-Adapter [180].      Fine-Tuning with LoRA        (§III-B5) Fine-grained: ReferDINO [104], VTG-LLM [58], VTimeLLM [63], ReVisionLLM [59], NumPro [185], InstFormer [57]. Coarse-grained: DiscoVLA [152], HierarQ [2].         Prompt Tuning        (§III-B6) Coarse-grained: ATP [9], ActionCLIP [171], X-CLIP [135], ViTiS [37], Q-ViD [133], VOP [65], CLIP-ViP [193].
@@
Figure 4: Taxonomy of image-to-video transfer learning paradigms and their corresponding methods. At top level, these methods can be grouped into two root paradigms: frozen features and adapted features, based on whether intermediate features extracted from the main information path of an ILFM remain totally unchanged or undergo adaptation. The frozen-features root paradigm includes paradigms of knowledge distillation, post-network tuning, and side-tuning, while the adapted-features root paradigm encompasses paradigms of full fine-tuning, partial tuning, fine-tuning with extra models, adapter, LoRA, and prompt tuning. For each specific transfer learning paradigm, we further divide its corresponding methods into fine-grained and coarse-grained types, reflecting different task requirements of the temporal and semantic alignment levels between visual and textual modalities.